Reports 1-1 of 1 Clear search Modify search
Detector Operation (Operations Report)
Gherardini - 8:44 Thursday 15 October 2020 (49664) Print this report
Operator Report - initial status
State of Virgo: Unlocked.

ITF Mode: Upgrading.

Quick Summary: IMC locked; SWEB & SNEB in-air; all suspensions loops closed with the exception of SR local controls; SR-BS Valve close; WE_UPS1 not working correctly; TCS CO2 lasers off; INJ_MAIN in pause, all the other automation nodes OFF.

Activities ongoing since this morning: mechanical works around SQZ area by LAPP team.
Images attached to this report
Comments to this report:
bertolini - 12:14 Thursday 15 October 2020 (49666) Print this report
I have recovered the SDB2 suspension by acting on the two fishing rods. I have also re-centered individually the two GAS filters that were both at the end of their mechanical range (one up one down).
Search Help